What is the molarity of 9.4g of sodium hydroxide mixed in 345mL of water?
1
Expert's answer
2017-04-07T10:28:07-0400
C_M (NaOH)=n(NaOH)/V(solution) =((m(NaOH))/(M(NaOH)))/(V(H_2 O))=(9.4g/(40g/mol))/0.345L=0.68 mol/L=0.68 M
Answer: 0.68 M.